# Causal Data Fusion for Panel Data without a Pre-Intervention Period

### v3

@@ A.4 Proof of Theorem 4
From Equations (16) and (17), we conclude that
$Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}=$ $~\varphi_{s}^{\top}(X_{1}-\sum_{i=2}^{J+1}w_{i}X_{i})$
$+\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{F_{1}^{R}-\sum_{i=2}^{J+1}w_{i}F_{i}^{R}\}$
$-\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}\}$
$+\tilde{\vartheta}_{s}^{\top}(\tilde{\mu}^{Y}_{1}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}_{i}^{Y})$
$+\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}$
$=$ $~-\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\{\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}\}$
$+\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}$
$+\tilde{\vartheta}_{s}^{\top}(\tilde{\mu}^{Y}_{1}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}_{i}^{Y}),$
where the last equality follows from Conditions (4)(4)-$(6)$.
Let
$R_{3s}=-\vartheta_{s}^{\top}({\theta^{R}}^{\top}\theta^{R})^{-1}{\theta^{R}}^{\top}\epsilon_{1}^{R}$
$R_{4s}=\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}$
$R_{5s}=\tilde{\vartheta}_{s}^{\top}(\tilde{\mu}_{1}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}_{i}),$
Then,
$Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}=R_{1s}+R_{2s}+R_{3s}+R_{4s}+R_{5s}.$
We note that the noise variables $\epsilon_{i,t}$ have mean zero, which gives us $\mathbb{E}[R_{3s}]=0$. $R_{4s}$ also has mean zero as the weights obtained from the reference domain are independent of mean-zero $\varepsilon_{i,s}$. Similar logic implies that $R_{4s}$ also has mean zero. However, $R_{1s}$ may not have mean zero as the weights are dependent on $\epsilon_{i,t}$. Similar concern holds for $R_{2s}$.
@@
and consequently, $b_{s}^{\top}\epsilon_{i}^{R}$ is sub-Gaussian with variance proxy $(\bar{\sigma}\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\,\frac{1}{\sqrt{T}})^{2}$. Consequently, applying Theorem 1.16 from (Rigollet and Hütter,, 2023), we have
@@
where $\mathbf{W}=\{w\in\mathbb{R}^{J};\sum w_{i}=1\}$.
@@
and consequently, $\tilde{b}_{s}^{\top}(\tilde{\mu}_{i}^{F}-\tilde{\mu}_{1}^{F})$ is sub-Gaussian with variance proxy $2(\tau\frac{\bar{\vartheta}\,\bar{\theta}\bar{\tilde{\theta}}}{\underline{\xi}})^{2}$. Consequently, applying Theorem 1.16 from (Rigollet and Hütter,, 2023), we have
@@
where $\mathbf{W}=\{w\in\mathbb{R}^{J};\sum w_{i}=1\}$.
Therefore, we can bound the bias of our proposed estimator as
$\left|\mathbb{E}[\hat{\psi}^{sc}]-\psi_{0}\right|$ $=\left|\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}\left[Y_{1,s}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}\right]-\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}\left[Y_{1,s}-Y_{1,s}^{(0)}\right]\right|$
@@
$\leq\frac{1}{S}\sum_{s=1}^{S}\left|\mathbb{E}\left[R_{1s}\right]+\mathbb{E}\left[R_{2s}\right]\right|$
$\leq\frac{1}{S}\sum_{s=1}^{S}\left|\mathbb{E}\left[R_{1s}\right]\right|+\left|\mathbb{E}\left[R_{2s}\right]\right|$
$\leq\sqrt{2}\bar{\vartheta}\bar{\theta}\bar{\sigma}\underline{\xi}^{-1}\sqrt{\frac{\log(2J)}{T}}+2\bar{\vartheta}\,\bar{\theta}\bar{\tilde{\theta}}\underline{\xi}^{-1}\sqrt{\log(2J)}.$

### v4

@@ Appendix A Proofs
Substituting into (16) yields
$Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}=$ $~\varphi_{s}^{\top}\Big(X_{1}-\sum_{i=2}^{J+1}w_{i}X_{i}\Big)$ (18)
$+\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}\Big(\bar{F}_{1}-\sum_{i=2}^{J+1}w_{i}\bar{F}_{i}\Big)$
$-\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}\phi^{R}\Big(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i}\Big)$
$-\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}\tilde{\theta}^{R}\Big(\tilde{\mu}_{1}^{F}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}_{i}^{F}\Big)$
$-\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}\Big(\epsilon_{1}^{R}-\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R}\Big)$
$+\tilde{\vartheta}_{s}^{\top}\Big(\tilde{\mu}_{1}^{Y}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}_{i}^{Y}\Big)+\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s}.$
Define the following terms:
$R_{1s}$ $:=\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}\sum_{i=2}^{J+1}w_{i}\epsilon_{i}^{R},$
$R_{2s}$ $:=-\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}\tilde{\theta}^{R}\Big(\tilde{\mu}_{1}^{F}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}_{i}^{F}\Big),$
$R_{3s}$ $:=-\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}\epsilon_{1}^{R},$
$R_{4s}$ $:=\varepsilon_{1,s}-\sum_{i=2}^{J+1}w_{i}\varepsilon_{i,s},$
$R_{5s}$ $:=\tilde{\vartheta}_{s}^{\top}\Big(\tilde{\mu}_{1}^{Y}-\sum_{i=2}^{J+1}w_{i}\tilde{\mu}_{i}^{Y}\Big),$
$R_{6s}$ $:=\varphi_{s}^{\top}\Big(X_{1}-\sum_{i=2}^{J+1}w_{i}X_{i}\Big),$
$R_{7s}$ $:=\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}\Big(\bar{F}_{1}-\sum_{i=2}^{J+1}w_{i}\bar{F}_{i}\Big),$
$R_{8s}$ $:=\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}\phi^{R}\Big(Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i}\Big).$
Then (18) can be written as
$Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}=R_{6s}+R_{7s}-R_{8s}+R_{1s}+R_{2s}+R_{3s}+R_{4s}+R_{5s}.$
By Assumption 5, $\mathbb{E}[R_{3s}]=0$. Moreover, since $\{\varepsilon_{i,s}\}_{i,s}$ and $\{\tilde{\mu}_{i}^{Y}\}_{i}$ are independent of $\mathcal{G}$ and have mean zero, and $w$ is $\mathcal{G}$-measurable, we have $\mathbb{E}[R_{4s}]=\mathbb{E}[R_{5s}]=0$. Therefore, by the triangle inequality,
$\left|\mathbb{E}\left[Y_{1,s}^{(0)}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}\right]\right|\leq|\mathbb{E}[R_{6s}]|+|\mathbb{E}[R_{7s}]|+|\mathbb{E}[R_{8s}]|+|\mathbb{E}[R_{1s}]|+|\mathbb{E}[R_{2s}]|.$ (19)
Bounding $|\mathbb{E}[R_{6s}]|,|\mathbb{E}[R_{7s}]|,|\mathbb{E}[R_{8s}]|$ using conditions (4)–(6) .
First, by Cauchy–Schwarz and $\|\varphi_{s}\|_{2}\leq\bar{\varphi}$,
$|\mathbb{E}[R_{6s}]|\leq\|\varphi_{s}\|_{2}\left\|X_{1}-\sum_{i=2}^{J+1}w_{i}X_{i}\right\|_{2}\leq\bar{\varphi}\sqrt{d_{t}}\,c,$
where the last step used condition (6)
Let $b_{s}:=\theta^{R}(\theta^{R^{\top}}\theta^{R})^{-1}\vartheta_{s}\in\mathbb{R}^{T}$. Then $\vartheta_{s}^{\top}(\theta^{R^{\top}}\theta^{R})^{-1}\theta^{R^{\top}}=b_{s}^{\top}$ and
$|\mathbb{E}[R_{7s}]|\leq\|b_{s}\|_{2}\left\|\bar{F}_{1}-\sum_{i=2}^{J+1}w_{i}\bar{F}_{i}\right\|_{2}.$
As in the standard argument,
$\|b_{s}\|_{2}\leq\|\theta^{R}\|_{\mathrm{op}}\;\|(\theta^{R^{\top}}\theta^{R})^{-1}\|_{\mathrm{op}}\;\|\vartheta_{s}\|_{2}\leq\frac{\sqrt{T}\,\bar{\theta}}{T\underline{\xi}}\,\bar{\vartheta}=\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\,\frac{1}{\sqrt{T}}.$
Combining with condition (4) i.e., $\|\bar{F}_{1}-\sum_{i}w_{i}\bar{F}_{i}\|_{2}\leq\sqrt{T}\,c$, gives
$|\mathbb{E}[R_{7s}]|\leq\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\,c.$
Similarly,
$|\mathbb{E}[R_{8s}]|\leq\|b_{s}^{\top}\phi^{R}\|_{2}\left\|Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i}\right\|_{2}\leq\|b_{s}\|_{2}\,\|\phi^{R}\|_{\mathrm{op}}\,\sqrt{d_{r}}\,c,$
using condition (5). Finally, $\|\phi^{R}\|_{\mathrm{op}}\leq\sqrt{\sum_{t=1}^{T}\|\phi_{t}\|_{2}^{2}}\leq\sqrt{T}\,\bar{\phi}$, hence
$|\mathbb{E}[R_{8s}]|\leq\left(\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\,\frac{1}{\sqrt{T}}\right)\left(\sqrt{T}\,\bar{\phi}\right)\sqrt{d_{r}}\,c=\frac{\bar{\vartheta}\,\bar{\theta}\,\bar{\phi}}{\underline{\xi}}\,\sqrt{d_{r}}\,c.$
Bounding $|\mathbb{E}[R_{1s}]|$ and $|\mathbb{E}[R_{2s}]|$.
@@
and consequently, $b_{s}^{\top}\epsilon_{i}^{R}$ is sub-Gaussian with variance proxy $(\bar{\sigma}\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\,\frac{1}{\sqrt{T}})^{2}$. Consequently, applying Theorem 1.16 from (Rigollet and Hütter, 2023), we have
@@
where $\mathbf{W}$ is the simplex
$\mathbf{W}:=\left\{w\in\mathbb{R}^{J}:~w_{i}\geq 0,\ \sum_{i=2}^{J+1}w_{i}=1\right\}.$
@@
and consequently, $\tilde{b}_{s}^{\top}(\tilde{\mu}_{i}^{F}-\tilde{\mu}_{1}^{F})$ is sub-Gaussian with variance proxy $2(\tau\frac{\bar{\vartheta}\,\bar{\theta}\bar{\tilde{\theta}}}{\underline{\xi}})^{2}$. Consequently, applying Theorem 1.16 from (Rigollet and Hütter, 2023), we have
@@
where $\mathbf{W}$ is the simplex
$\mathbf{W}:=\left\{w\in\mathbb{R}^{J}:~w_{i}\geq 0,\ \sum_{i=2}^{J+1}w_{i}=1\right\}.$
Then
$|\mathbb{E}[R_{1s}]|\leq\sqrt{2}\,\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\,\bar{\sigma}\sqrt{\frac{\log(2J)}{T}},\qquad|\mathbb{E}[R_{2s}]|\leq 2\,\frac{\bar{\vartheta}\,\bar{\theta}\,\bar{\tilde{\theta}}}{\underline{\xi}}\,\tau\,\sqrt{\log(2J)}.$
Conclusion.
Combining these bounds with (19) and averaging over $s=1,\ldots,S$,
$\big|\mathbb{E}[\hat{\psi}^{sc}]-\psi_{0}\big|$ $=\left|\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}\left[Y_{1,s}-\sum_{i=2}^{J+1}w_{i}Y_{i,s}\right]-\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}\left[Y_{1,s}-Y_{1,s}^{(0)}\right]\right|$
@@ Conclusion.
$\leq\Bigg(\bar{\varphi}\sqrt{d_{t}}+\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}+\frac{\bar{\vartheta}\,\bar{\theta}\,\bar{\phi}}{\underline{\xi}}\sqrt{d_{r}}\Bigg)c+\sqrt{2}\,\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\,\bar{\sigma}\sqrt{\frac{\log(2J)}{T}}+2\,\frac{\bar{\vartheta}\,\bar{\theta}\,\bar{\tilde{\theta}}}{\underline{\xi}}\,\tau\,\sqrt{\log(2J)}.$
This proves the theorem.